OpenStudy (anonymous):

limit x->5 ((x^3-6*x^2+25)/(x-5) the answer of this question is 15 which we can determine by putting such values of of x which are close to 5 (but not exactly 5), for example first I started with these values 4.8,4.9,4.99 & -4.8,-4.9,-4.99. It's kind of a lengthy process, is there any quick process for such problem?

OpenStudy (anonymous):

limit's are not just about substituting... we have to do some algebra first (if we can !) you have denominator x-5 so substituting would be like 5-5=0 and we don't want that to happen.... so note that we can factor numerator ... and then we simplify... here's how it goes ! \[\LARGE \lim_{x\to 5}\frac{x^3-6x^2+25}{x-5}=\lim_{x\to5}\frac{(x-5)(x^2-x-5)}{x-5}\] \[\LARGE \lim_{x\to5}(x^2-x-5)=5^2-5-5=25-10=15\] Hope this helps :)

OpenStudy (anonymous):

what if we can't simplify the equation or if the equation is already simplified

OpenStudy (anonymous):

post an example.. let's see :(

OpenStudy (anonymous):

it's OK no problem thanks anyway

OpenStudy (anonymous):

ok... as you wish ;) Glad to help, you're welcome.